OpenStudy (ars22):

how to factor -x^2 - 2x + 15

OpenStudy (oaktree):

First divide everything by -1.

OpenStudy (oaktree):

You get x^2+2x-15

OpenStudy (oaktree):

It factors then as (x+5)(x-3) and you're done.

OpenStudy (anonymous):

(-x-5)(x-3)

OpenStudy (anonymous):

(-x-5)(x-3)

OpenStudy (oaktree):

No, you need to factor a (-1) out of your (-x-5) term. Your leading term cannot be negative.

OpenStudy (anonymous):

Well, then you'd better include it in your final answer (-1)(x+5)(x-3).

OpenStudy (oaktree):

No, because you can't factor originally with a negative x^2.

OpenStudy (oaktree):

So that factoring is wrong.

OpenStudy (ars22):

-(x+5)(x-3)

OpenStudy (ars22):

-(x+5)(x-3) this is the right answer i think but how

OpenStudy (anonymous):

Well you can begin like OakTree and pull out the -1: -(x^2+2x-15); then you are looking for something like (x+a)(x+b) but what are a and b? Consider what that looks like when multiplied out: x^2+ax+bx+ab. The x^2 is right. ax+bx has to equal 2x; therefor a+b=2 ab must equal -15 So consider the factor pairs of -15 that could be a and b: 1,-15; -1,15; 3,-5; -3,5 Only the last pair adds to 2, so it must be -(x-3)(x+5)

OpenStudy (ars22):

thanks

OpenStudy (anonymous):

-x^2 - 2x + 15 = -(x^2+2x-15) = - (x-5)(x+3)

OpenStudy (ars22):

no, both x's would be negative

OpenStudy (ars22):

no negative outside

OpenStudy (anonymous):

it's ok to have negative outside too, both ways work.